solve for n 4m+5n=25
Answer by LagrangeSon678(5)

We have 4m+5b=25
To solve for n we have to isolate it on one side of the equation.
Lets begin by subtracting 4m from both side, this results in: 5n=25-4m.
The final step is to divide both sides by 5.
Which gives us: n= (25-4m)/(5)
That is our answer.